# Experiences from Using LLMs for Repository Mining Studies in Empirical Software Engineering

### v1

@@ III Using LLMs for ESE: Methodological Insights
Figure 2 shows the proposed framework PRIMES, based on experiences from  [9, 10]. The PRIMES has four stages, mapping to the four subsections below and highlighted with numbers from to . Each stage consists of several steps, which are highlighted in italic in the text below, and map to small boxes in dark blue in Figure 2. These steps can be seen as a preliminary checklist when designing repository mining studies using LLMs for automated data collection.
@@ III-C Evaluation among multiple LLMs ()
Benchmark: Establishing a benchmark is crucial for fairly evaluating a prompt’s performance across multiple LLMs. Improving a prompt based on a single LLM can introduce biases from that model. To avoid this, involving multiple LLMs in the prompt creation process ensures a balanced evaluation to also assess the generalizability of the prompt. It is essential to establish a set of metrics to compare performance and objectively align assessments with specific goals.
@@ III-D Output Validation ()
The output validation is essential for ensuring the data collection and analysis of LLM results, facilitating replication, and enabling future studies to build upon the work. Several handling issues may arise when evaluating model outputs. When validation uncovers issues such as hallucinations, duplications, or formatting errors, systematically address them.
@@ IV Discussions and Outlook
To address these challenges, we implemented mitigation strategies, including rigorous validation procedures using expert-based oracles to evaluate LLM outputs and reduce inaccuracies. We assessed output reliability with statistical measures like Cohen’s kappa and compared multiple LLMs based on metrics such as classification accuracy, interpretability, and cost-effectiveness to find the most suitable model. We acknowledge that model biases and the black-box nature of LLMs still require further research and attention. Moreover, challenges such as ensuring data privacy and addressing ethical concerns when using proprietary LLMs necessitate careful consideration. The environmental impact of training and deploying large-scale LLMs due to high computational demands is another area that merits attention. Additionally, integrating LLMs effectively into existing SE workflows presents practical difficulties. Future research should explore these issues to fully harness the potential of LLMs in ESE.
@@
Extension and Adaptability of the Proposed Framework: While our study focused on particular aspects of software repository mining, the methods and insights gained can be adapted to various contexts. Future research could investigate the relevance of our results across different software domains to assess the generalizability of our framework PRIMES. Additionally, incorporating feedback from subsequent studies and user experiences will enable iterative improvements, enhancing the framework’s adaptability to diverse analytical scenarios, such as code quality assessment and automated testing. One potential direction includes the creation and extension of a metrics catalog to benchmark multiple LLMs. Exploring the integration of emerging technologies can further enrich the framework’s capabilities, making it a robust resource for researchers and practitioners.

### v2

# A Framework for Using LLMs for Repository Mining Studies in Empirical Software Engineering

@@ III Using LLMs for ESE: Methodological Insights
Figure 2 shows the proposed framework PRIMES, based on experiences from  [9, 10]. PRIMES has four stages, mapping to the four subsections below and highlighted with numbers from to . Each stage consists of several steps, which are highlighted in italic in the text below, and map to small boxes in dark blue in Figure 2. These steps can be seen as a preliminary checklist when designing repository mining studies using LLMs for automated data collection.
@@ III-C Evaluation among multiple LLMs ()
Benchmark: Establishing a benchmark is crucial for fairly evaluating a prompt’s performance across multiple LLMs. A poorly designed benchmark can lead to overconfidence in the LLM outputs, potentially creating unrealistic scenarios that fail to reveal limitations.
Design a benchmark with tests to assess generalization, including false positives to test accuracy and robustness, and ensure task diversity to assess adaptability. This should be guided by a comprehensive set of metrics to compare performance and align assessments with specific goals.
@@ III-D Output Validation ()
The output validation is essential for ensuring the data collection and analysis of LLM results, facilitating replication, and enabling future studies to build upon the work.
Several handling issues may arise when evaluating model outputs. When validation uncovers issues such as hallucinations, duplications, or formatting errors, these have to be systematically addressed.
@@ IV Discussions and Outlook
To address these challenges, we implemented mitigation strategies, including rigorous validation procedures using expert-based oracles to evaluate LLM outputs and reduce inaccuracies. We assessed output reliability with statistical measures like Cohen’s kappa and compared multiple LLMs based on metrics such as classification accuracy, interpretability, and cost-effectiveness to find the most suitable model. We acknowledge that model biases and the black-box nature of LLMs still require further research and attention. Moreover, challenges such as ensuring data privacy and addressing ethical concerns when using proprietary LLMs necessitate careful consideration. Another critical but often overlooked challenge is the environmental cost impact of training and deploying large-scale LLMs. Incorporating this dimension into the assessment framework ensures that trade-offs are aligned with sustainable AI principles. Additionally, integrating LLMs effectively into existing SE workflows presents practical difficulties. Future research should explore these issues to fully harness the potential of LLMs in ESE.
@@
Extension and Adaptability of the Proposed Framework: While our study focused on particular aspects of software repository mining, the methods and insights gained can be adapted to various contexts. Future research could investigate the relevance of our results across different software domains to assess the generalizability of our framework PRIMES. Additionally, incorporating feedback from subsequent studies and user experiences will enable iterative improvements, enhancing the framework’s adaptability to diverse analytical scenarios, such as code quality assessment and automated testing. One potential direction includes the creation and extension of a metrics catalog to benchmark multiple LLMs and consider their trade-off. Exploring the integration of emerging technologies can further enrich the framework’s capabilities, making it a robust resource for researchers and practitioners.